# Variational Sparse Paired Autoencoders (vsPAIR) for Inverse Problems and Uncertainty Quantification

### v1

@@ 4 Numerical Results
We evaluate vsPAIR on two inverse problems: blind inpainting on MNIST and computed tomography (CT) on the LoDoPaB-CT dataset [36]. The first is a controlled setting where we examine reconstruction quality, sparsity of the learned representations, and the interpretability of uncertainty estimates. The second inverse problem concerns CT, and whether these properties hold on a larger-scale problem with more complex image structure and realistic noise. To ensure reproducibility, elementary code will be made available on Github upon initial acceptance. Details regarding hyperparameter tuning can be found in Appendix A.
@@ 4.1 Blind Inpainting on MNIST
All models use the same convolutional neural network architecture and were trained for 100 epochs. Architecture details are provided in Section B.1. Moreover, all decoders are deterministic.
We organize the MNIST evaluation around three questions. First, how does reconstruction quality change as we introduce variational structure and sparsity? Second, do the learned latent representations capture interpretable features, and can individual dimensions be associated with semantically meaningful variation in the reconstructions? Third, does the combination of paired structure and sparsity yield structured uncertainty estimates — that is, can the model not only identify where it is uncertain but also learn latent factors that selectively control the uncertain regions? We address these questions through a combination of quantitative metrics and qualitative analyses across the four methods.
@@
Table 3 reports reconstruction error and sparsity metrics across the test set. We tune the sparse models to yield a number of active dimensions comparable to the 32-dimensional dense baselines. This sparsity level is tunable via the beta distribution hyperprior parameters $(\alpha_{0},\beta_{0})$, which is discussed further in Appendix A.
@@
Table 4: Consistency of active latent dimensions across the MNIST test set. For each image, we count dimensions active in all 30 stochastic samples. Consistently active dimensions form the basis for the localization analysis in Figure 8, as only stable dimensions permit reliable attribution of learned factors. vPAIR uses all dimensions by construction, vsPAIR maintains a moderate stable subset, and sVAE shows high variability due to its stochastic gating mechanism.
@@ 4.2 Computed Tomography
For our experiments, we train and test on a subset of the dataset comprising 3,000 training images and 800 test images, randomly selected from the full collection. Both the original images and sinograms are cropped and down-sampled by a factor of four, yielding a QoI of size $88\times 88=7,744$ and observations of size $248\times 128=31,744$. The vsPAIR model employs CNN-based encoder–decoder architectures, with latent dimensions of $8,192$ for the QoI encoder and $2,048$ for the observation encoder.
The first experiment establishes our method as a competitive end-to-end inversion approach. We simulate reconstruction at a variety of noise levels. The lowest level only includes simulated Poisson noise present in the LoDoPaB-CT dataset. We further generate datasets with $2.5,5,10$ and $20\%$ Gaussian white noise added to the sinograms in order to simulate high-noise settings. Five separate vsPAIR models were trained for 100 epochs on each noise level, for details on hyperparameters and training, see Section B.2.
Example results for each noise level tested are shown in Figure 9. Here, we also show reconstructions using Filtered Back Projection (FBP)[29] as a baseline, and compare to a more modern data-driven method, Learned Primal-Dual (LPD) [2] which performs a standard unrolled proximal primal-dual optimization method, replacing the proximal operators with convolutional neural networks.
@@
From Figure 9, it is evident from the results that vsPAIR offers competitive reconstruction performance to both LPD and FBP. As expected, standard iterative approaches (FBP) fail to provide meaningful reconstructions as the noise level increases, while the two data-driven approaches, LPD and vsPAIR are capable of maintaining meaningful reconstructions. In particular, vsPAIR outperforms LPD for high-noise applications. It should be noted that these methods are often paired with further post-processing methods for use in practice [36], and the application of vsPAIR in these broader frameworks remains to be studied.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 10: LPD and vsPAIR reconstructions trained in the variable noise setting. vsPAIR plots are overlaid with UQ estimation, generated by drawing 1,000 examples from the latent distribution, and computing the pixel-wise variance between samples. All images are normalized to the same $[0,1]$ color-scale.
For the remaining results, we focus on a variable noise level, ranging from $0\%$ to $20\%$ Gaussian white noise in addition to the simulated Poisson noise, and aim to illustrate the UQ advantage of the vsPAIR framework. Figure 10 shows that both LPD and vsPAIR remain competitive in the case of variable noise. The latent representations of the CT images had 3,150 nnz elements, on average. Moreover, implementation using variational autoencoders allows for UQ inaccessible for deterministic approaches. From Figure 10, we observe that the pixel-wise variance for each reconstruction is small, consistent with the over-confidence often exhibited by VAEs [15]. Although small, the relative value of the variances offer useful insight into the generated images. The model is often highly uncertain in regions of fine details and the edges of different regions, which are often incorrect or obscured in the reconstruction. Thus, in addition to offering competitive surrogate inversion, the vsPAIR framework is able to clearly distinguish regions of high uncertainty through sampling of the latent variable $\hat{z}_{x}$.
@@
Using the same model as in the previous experiment, we analyze latent variability by perturbing the mean latent representation. We select latent features with the highest variance and activation probability exceeding $95\%$ ($\omega_{i}\geq 0.95$), and visualize the resulting reconstructions as each dimension is varied by multiples of its variance; see Figure 11. Extreme variations are included to emphasize the semantic effect of individual latent dimensions.
Figure 11: These figures show the effect of varying the latent indices with the top 3 highest variances that also have a higher than $95\%$ chance of being active. Here, two different reconstructions are shown where the selected latent indices are varied within 2 standard distributions of their mean (left) paired with a plot showing extreme perturbation of these values (up to 200 standard deviations) to further illustrate the effects of the index on reconstruction (right). The rightmost column of each subplot shows the pixel-wise variance across the perturbations.
In the first and second row of the first image, we observe that extreme perturbations of the indices with the first and second highest variance correspond to changes in the outline and edges of the scan, as well as fine detail in the inner dark cavity. The third index appears to focus on the boundary of the inner cavity and the bone structure to the right of the image. This indicates that the index that the model is most uncertain about has a direct influence on these details, which is consistent with the findings illustrated in Figure 10. A more interesting case concerns the small black hole in the second example. Here, we observe that varying the index ranked second in uncertainty has a direct effect on the size of the hole. From this analysis, we can conclude that the corresponding index encodes some information about the hole, and from model parameters, we can conclude that it is uncertain of the size and boundaries of it. It should be noted that for images uncharacteristic of the dataset, indices with high uncertainty often did not have an interpretable effect on the reconstruction.
@@
A key advantage of the vsPAIR approach over LPD and FBP is that vsPAIR does not require access to the forward operator. This certainly comes at the cost of model complexity, as the vsPAIR approaches require significantly more parameters and training time: the entire vsPAIR framework used for this example had approximately $1.1$ billion parameters, while LPD only required approximately $251$ thousand. Of course, various existing approaches could be leveraged to reduce the size of the vsPAIR networks [6, 35] and should be implemented for improved practicality. Additionally, a limitation of this work is that only CNN architectures are used for the models, and future work will further investigate the use of other networks.
@@ 5 Discussion and Conclusion
The methodology was validated both numerically and with theoretical motivation. From a theoretical standpoint, we established conditions under which the latent distribution-mapping training objective is well posed, and showed that for linear inverse problems with Gaussian noise the framework can learn a mapping to the conditional distribution of the QoI latent representation. These results rely on restrictive assumptions on the encoder and serve primarily to motivate the architecture rather than guarantee performance in general settings: future work expands upon this analysis and further investigate the theoretical properties of paired VAE frameworks. The practical behavior of vsPAIR on nonlinear problems is assessed empirically.
@@
There are many avenues for future work. Reducing model size through compression techniques or more efficient architectures would improve scalability. Additional constraints on the latent space may further enhance semantic interpretability. The framework could also be extended to other inverse problems where interpretable uncertainty is valuable, particularly those with inherently low-dimensional structure where sparse representations may naturally align with true factors of variation.
@@ Acknowledgments.
Appendix A Hyperparameter Selection
@@
It should be noted that the ELBOs for both the sVAE and VAE are typically broken up into a reconstruction and KL term as shown in Equation 3. It is standard to assign a hyperparameter weight to the term $\rm{D}_{\rm KL}$ to enforce the stability of training, which is the defining characteristic of a $\beta$-VAE. The effects of this hyperparameter on VAE training are well studied in other works [9, 46, 38]. We denote the weight of the KL term in the QoI and observation ELBO as $\gamma_{x}$ and $\gamma_{y}$, respectively, and found that weighting the KL terms of both sVAE and VAE by $\gamma_{x}=1.0$ and $\gamma_{y}=0.1$ resulted in more stable training and higher fidelity reconstructions without significantly compromising the variance of the reconstructions.
Appendix B Model Architectures and Training Details
B.1 MNIST Example

### v2

@@ 4 Numerical Results
We evaluate vsPAIR on two inverse problems: blind inpainting on MNIST and computed tomography (CT) on the LoDoPaB-CT dataset [40]. The first is a controlled setting where we examine reconstruction quality, sparsity of the learned representations, and the interpretability of uncertainty estimates. The second inverse problem concerns CT, and whether these properties hold on a larger-scale problem with more complex image structure and realistic noise. To ensure reproducibility, elementary code will be made available on Github upon initial acceptance. Details regarding hyperparameter tuning can be found in Appendix B.
@@ 4.1 Blind Inpainting on MNIST
All models use the same convolutional neural network architecture and were trained for 100 epochs. Architecture details are provided in Section D.1. Moreover, all decoders are deterministic.
We organize the MNIST evaluation around three questions. First, how does reconstruction quality change as we introduce variational structure and sparsity? Second, do the learned latent representations capture interpretable features, and can individual dimensions be associated with semantically meaningful variation in the reconstructions? Third, does the combination of paired structure and sparsity yield structured uncertainty estimates–that is, can the model not only identify where it is uncertain but also learn latent factors that selectively control the uncertain regions? We address these questions through a combination of quantitative metrics and qualitative analyses across the four methods.
@@
Table 3 reports reconstruction error and sparsity metrics across the test set. We tune the sparse models to yield a number of active dimensions comparable to the 32-dimensional dense baselines. This sparsity level is tunable via the beta distribution hyperprior parameters $(\alpha_{0},\beta_{0})$, which is discussed further in Appendix B.
@@
Table 4: Consistency of active latent dimensions across the MNIST test set. For each image, we count dimensions active in all 30 stochastic samples. Consistently active dimensions form the basis for the localization analysis in Figure 8, as only stable dimensions permit reliable attribution of learned factors. vPAIR uses all dimensions by construction, vsPAIR maintains a moderate stable subset, and sVAE shows high variability.
@@ 4.2 Computed Tomography
For our experiments, we train and test on a subset of the dataset comprising 3,000 training images and 800 test images, randomly selected from the full collection. Both the original images and sinograms are cropped and down-sampled by a factor of four, yielding a QoI of size $88\times 88=7{,}744$ and observations of size $248\times 128=31{,}744$. The vsPAIR model employs CNN-based encoder–decoder architectures, with latent dimensions of 8,1928{,}192 for the QoI encoder and 2,0482{,}048 for the observation encoder.
The first experiment establishes our method as a competitive end-to-end inversion approach. We simulate reconstruction at a variety of noise levels. The lowest level only includes simulated Poisson noise present in the LoDoPaB-CT dataset. We further generate datasets with $2.5,5,10$ and $20\%$ Gaussian white noise added to the sinograms in order to simulate high-noise settings. Five separate vsPAIR models were trained for 100 epochs on each noise level. Hyperparameters were configured for high-fidelity reconstructions–for details, see Section D.2.
Example results for each noise level tested are shown in Figure 9. Here, we also show reconstructions using Filtered Back Projection (FBP)[32] as a baseline, and compare to a more modern data-driven method, Learned Primal-Dual (LPD) [2, 20] which performs a standard unrolled proximal primal-dual optimization method, replacing the proximal operators with convolutional neural networks.
@@
From Figure 9, it is evident from the results that vsPAIR offers competitive reconstruction performance to both LPD and FBP. As expected, standard iterative approaches (FBP) fail to provide meaningful reconstructions as the noise level increases, while the two data-driven approaches, LPD and vsPAIR are capable of maintaining suitable reconstructions. In particular, vsPAIR outperforms LPD for high-noise applications. It should be noted that these methods are often paired with further post-processing methods for use in practice [40], and the application of vsPAIR in these broader frameworks remains to be studied.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 10: vsPAIR reconstructions trained in the variable noise setting. The rightmost column shows the mean overlaid with UQ estimation, generated by drawing 1,000 examples from the latent distribution, and computing the pixel-wise variance between samples. All images are normalized to the same $[0,1]$ color-scale.
For the remaining results, we focus on a variable noise level, ranging from $0\%$ to $20\%$ Gaussian white noise in addition to the simulated Poisson noise, and aim to illustrate the UQ advantage of the vsPAIR framework. Figure 10 shows that vsPAIR provides decent reconstructions in the case of variable noise. The latent representations of the CT images had 7,041 nnz elements, on average. Moreover, implementation using variational autoencoders allows for UQ inaccessible for deterministic approaches. From Figure 10, we observe that the pixel-wise variance for each reconstruction is small, consistent with the over-confidence often exhibited by VAEs [16]. Although small, the relative value of the variances offer useful insight into the generated images. The model is often highly uncertain in regions of fine details and the edges of different regions, which are often incorrect or obscured in the reconstruction. Thus, in addition to offering competitive surrogate inversion, the vsPAIR framework is able to distinguish regions of high uncertainty through sampling of the latent variable $\hat{z}_{x}$.
@@
Using the same approach as in the previous experiment, we analyze latent variability by perturbing the mean latent representation. We select latent features with the highest variance and activation probability exceeding $95\%$ ($\omega_{i}\geq 0.95$), and visualize the resulting reconstructions as each dimension is varied by multiples of its variance. We repeat this approach for the figures with the lowest variance above the $95\%$ threshold; see Figure 11.
Figure 11: These figures show the effect of varying the latent indices with the top 3 highest (left) and top 3 lowest (right) variances that also have a higher than $95\%$ chance of being active. Here, two different reconstructions are shown where the selected latent indices are varied 1 standard distributions from their mean. The rightmost column of each subplot shows the pixel-wise variance across the perturbations of the index.
The results in Figure 10 align with those of Figure 11 in that the model exhibits high uncertainty in fine details of the image. Between both images in Figure 11, we notice a consistent trend: the features with the lowest variance seemingly correspond to large dominant features of the image, while the images with the highest variance seemingly correspond to fine details. For example, in the first image (top), the perturbation of low variance features affects the outline and edges of the inner dark cavity, while perturbation of the high-variance features corresponds to the details inside the cavities, as well as the pixel values of the white structure in the center-right. This trend is evident in the second image second image (bottom), where the low-variance features indicate confidence in the structure of the image–i.e. the hole as as well as the edges. Perturbation of the high-variance features also reveals a change in the pixel values surrounding the hole. From this analysis, it appears that the model is confident in the hole existing, but uncertain about the exact size or shape. It should be noted that for images uncharacteristic of the dataset, specific indices did not have an interpretable effect on the reconstruction. Moreover, the latent indices that seemingly correspond to image features is heavily dependent on each specific image and the added noise.
@@
A key advantage of the vsPAIR approach over LPD and FBP is that vsPAIR does not require access to the forward operator. This certainly comes at the cost of model complexity, as the vsPAIR approaches require significantly more parameters and training time: the entire vsPAIR framework used for this example had approximately $1.1$ billion parameters, while LPD only required approximately $251$ thousand. Of course, various existing approaches could be leveraged to reduce the size of the vsPAIR networks [5, 39] and should be implemented for improved practicality. Additionally, a limitation of this work is that only CNN architectures are used for the models, and future work will further investigate the use of other networks and sVAE configurations.
@@ 5 Discussion and Conclusion
The methodology was validated both numerically and with theoretical motivation. From a theoretical standpoint, we established conditions under which the latent distribution-mapping training objective is learnable, and showed that for linear inverse problems with Gaussian noise the framework can learn a mapping to the conditional distribution of the QoI latent representation. These results rely on restrictive assumptions on the encoder and serve primarily to motivate the architecture rather than guarantee performance in general settings: future work expands upon this analysis and further investigate the theoretical properties of paired VAE frameworks. The practical behavior of vsPAIR is assessed empirically.
@@
There are many avenues for future work. Reducing model size through compression techniques or more efficient architectures would improve scalability and more advanced implementations of sVAEs can be analyzed. Additional constraints on the latent space may further enhance semantic interpretability. The framework could also be extended to other inverse problems where interpretable uncertainty is valuable, particularly those with inherently low-dimensional structure where sparse representations may naturally align with true factors of variation.
@@ Acknowledgments.
Appendix A sVAE Implementation Details
Beta Hyperprior for Adaptive Sparsity.
Introducing sparsity into an sVAE is not trivial and a sparsity level, $(1-\rho)$, for the prior must be specified. This can be incorporated index-wise, or fixed for all elements. [62] originally leveraged a classifier-based approach that builds off of [60], where a classifier is used to pick a specific prior for each input, based on trained pseudo-inputs. This motivates structure in the latent representations.
For simplicity, we fix the prior level of sparsity $\rho$ for all inputs and indices, instead analyzing this problem in a Bayesian sense. We assume that the distribution of $\rho$ is described by some underlying hyperprior distribution. We treat $\rho$ as a learnable parameter regularized by a Beta hyperprior
$\rho\sim\mathrm{Beta}(\alpha_{0},\beta_{0}).$ (20)
The use of a Beta prior is common in sparse coding–see [48, 68]. In particular, our approach is similar to [53], yet different in that the previous work prescribes a Beta prior to the mixture weights of the spike-and-slab encoder, while we enforce it on the sparsity level of the prior. We consider the negative log-density as a training objective dependent on $\rho$,
$\mathcal{L}_{\rho}(\rho\,;\,\alpha_{0},\beta_{0})=(\alpha_{0}-1)\log\rho+(\beta_{0}-1)\log(1-\rho).$ (21)
Here, terms constant in $\rho$ are dropped. The likelihood term seeks to push $\rho$ towards a probable value under the prescribed hyperprior, but is competing with the sVAE loss term. This establishes a tradeoff between sVAE performance and sparsity, for which optimization finds a balance. Since $\rho$ represents the prior probability of a coordinate being active, larger $\rho$ corresponds more dense reconstructions. Setting $\alpha_{0}>\beta_{0}$ yields a left-skewed beta density that concentrates mass toward higher values of $\rho$, encouraging dense latent representations. Conversely, sparsity is favored when $\alpha_{0}<\beta_{0}$. Numerical results confirm that this approach aligns with these expectations–see Appendix B.
Bernoulli Relaxation.
Training generative models under sparsity constraints is a well-studied field [43, 6]. A central challenge in training sVAEs arises from the non-differentiable Bernoulli sampling procedure: the loss involves quantities computed from draws of Bernoulli trials. The approach in [62], which builds off of [45] and [54], addresses this issue via a differentiable relaxation of Bernoulli sampling based on a scaled sigmoid and we follow the notation in [62]. In particular, the latent encoding is computed element-wise as
$z_{i}=T(\eta_{i}-1+\omega_{i})\;\cdot\;(\mu_{i}+\sigma_{i}\epsilon)$ (22)
To enforce true sparsity, we leverage a straight-through gradient estimator [6] similar to the approach used in [21]. Here, the gradient of the Bernoulli relaxation is used in backpropagation, while thresholding is used to enact exact zeros in the forward pass. This is implemented as follows:
  1. 1.
    Sample uniform noise $\eta_{i}\sim\mathrm{Uniform}(0,1)$.
  2. 2.
    Compute a relaxation of the thresholding via the differentiable step function approximation: $s_{i}=\sigma\bigl(c(\eta_{i}-1+\omega_{i})\bigr)\in(0,1)$, where $c>0$ is a temperature parameter ([62, 45]).
  3. 3.
    Threshold to obtain exact zeros: $\hat{s}_{i}=\mathbf{1}_{\{s_{i}>0.5\}}(s_{i})\in\{0,1\}$.
  4. 4.
    Straight-through gradient estimation: construct a mask, ${\rm mask}_{i}$, using the thresholded values, but use the gradient of $s_{i}$ for backpropagation [6, 21, 63].
  5. 5.
    Apply the standard reparameterization ${z}_{i}^{*}=\mu_{i}+\sigma_{i}\epsilon_{i}$ with $\epsilon_{i}\sim\mathcal{N}(\epsilon\mid 0,1)$, then mask: $z_{i}=\mathrm{mask}_{i}\cdot z_{i}^{*}$ [35].
Appendix B Hyperparameter Selection
@@
It should be noted that the ELBOs for both the sVAE and VAE are typically broken up into a reconstruction and KL term as shown in Equation 3. It is standard to assign a hyperparameter weight to the term $\rm{D}_{\rm KL}$ to enforce the stability of training, which is the defining characteristic of a $\beta$-VAE. The effects of this hyperparameter on VAE training are well studied in other works [10, 51, 42]. We denote the weight of the KL term in the QoI and observation ELBO as $\gamma_{x}$ and $\gamma_{y}$, respectively, and found that weighting the KL terms of both sVAE and VAE by $\gamma_{x}=1.0$ and $\gamma_{y}=0.1$ resulted in more stable training and higher fidelity reconstructions without significantly compromising the variance of the reconstructions. Implicitly, $\gamma_{x}$ also controls the sparsity level since it reflects the strength of the regularization that pushes the encoder to the prior. The results in the CT study used a different configuration with lower $\gamma_{x}$ and $\gamma_{y}$ weights to promote reconstruction accuracy.
Appendix C Theoretical Results
C.1 Proof of Theorem 1
The expression for the conditional mean follows directly from the tower property of conditional expectation and the conditional independence assumption [33]:
$\mathbb{E}[Z_{x}\mid Y]=\mathbb{E}[\mathbb{E}[Z_{x}\mid X,Y]\mid Y]$ (23)
$=\mathbb{E}[\mathbb{E}[Z_{x}\mid X]\mid Y]$ (24)
$=\mathbb{E}[\mu_{x}(X)\mid Y]=B\mathbb{E}[X\mid Y]+c.$ (25)
The last line follows from plugging in the affine expression for $\mu_{x}(X)$ and the linearity of expectation.
The covariance expression follows directly from the law of total covariance and conditional independence [33]. From this result, we have
${\rm Cov}(Z_{x}\mid Y)=\mathbb{E}[{\rm Cov}(Z_{x}\mid X,Y)\mid Y]+{\rm Cov(}\mathbb{E}[(Z_{x}\mid X,Y)]\mid Y),$ (26)
which by the conditional independence gives
$\mathbb{E}({\rm Cov}(Z_{x}\mid X)\mid Y)+{\rm Cov}(\mathbb{E}[Z_{x}\mid X]\mid Y).$ (27)
Now, we substitute the encoder mean and covariance, which gives
$\mathbb{E}(\Sigma_{x}(X)\mid Y)+{\rm Cov}(\mu_{x}(X)\mid Y),$ (28)
and simplifies to
$\Sigma_{Z_{x}}+B{\rm Cov}(X\mid Y)B^{\top}$ (29)
since $\Sigma_{x}(x)$ is constant, and $\mu_{x}(x)$ is affine. This is clearly symmetric positive definite as it is the sum of symmetric positive definite and symmetric positive semi-definite matrices. Continuity follows from the first assumption. More general results can be obtained by enforcing more regularity in the parameter maps and in the conditional expectation and covariance.
C.2 Proof of Theorem 2
Let $Y=AX+\varepsilon$ with $\varepsilon$ and $p(x\mid z_{x})$ Gaussian distributed. Writing $p(y\mid z_{x})$ as the marginal over $x$, we obtain
$p(y\mid z_{x})=\int_{X}p(y,x\mid z_{x})dx$ (30)
$p(y\mid z_{x})=\int_{X}p(y\mid x,z_{x})p(x\mid z_{x})dx$ (31)
$p(y\mid z_{x})=\int_{X}p(y\mid x)p(x\mid z_{x})dx$ (32)
by our conditional independence assumption. The linearity of the forward map implies that $p(y\mid x)$ will be Gaussian [11]. Note that this matches the marginal likelihood computation for Bayesian linear regression with a Gaussian prior. This integral can be computed in closed form, and results in a Gaussian distribution [8, 46].
By Bayes Theorem,
$p(z_{x}\mid y)\propto p(y\mid z_{x})p(z_{x})$ (33)
which is Gaussian, under the prior assumption that $p(z_{x})$ is Gaussian as it constitutes the product of two Gaussian densities. The result in Theorem 2 is obtained by replacing $p(x\mid z_{x})$ by $q_{\theta_{x}}(x\mid z_{x})$.
Appendix D Model Architectures and Training Details
D.1 MNIST Example